40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 4133     Entry time: Tue Jan 11 11:39:45 2011
Author: steve 
Type: Update 
Category: VAC 
Subject: RGA is turned on 

Joe updated the rga procedure in the wiki and we turned on the filament. It will take a few days to reach thermal equilibrium.

TP2 dry fore pump was replaced at 910mTorr

Attachment 1: pd70d22.jpg  68 kB  | Hide | Hide all
pd70d22.jpg
ELOG V3.1.3-